# Security Analysis of Web Applications Based on Gruyere

### v1

@@ 1. Introduction
With the rapid advancement of internet technology, web systems have become a critical infrastructure for exchanging information and processing business. However, their complexity has also led to the emergence of numerous security vulnerabilities, making web security a core research focus in cybersecurity. Web security not only concerns data protection and privacy but also directly impacts business continuity and system reliability(chen2013dynamic, ). Therefore, conducting web security research holds significant theoretical and practical importance for mitigating malicious attack risks and enhancing system robustness. Gruyere is a web application platform developed by Google specifically for security research. It integrates multiple typical vulnerabilities, covering cross-site scripting (XSS), cross-site request forgery (CSRF), and access control flaws(douligeris2004ddos, ). The platform is easy to operate and closely mirrors real-world scenarios, providing researchers with an experimental environment for vulnerability analysis and verification while serving as an ideal tool for teaching and training. Research based on Gruyere can reveal shortcomings in existing protective mechanisms, drive the development of novel defense technologies, and provide references for the industry in vulnerability protection and risk control(zivanic2022network, ).
In recent years, scholars worldwide have made positive progress in the field of web security. However, deficiencies remain in vulnerability detection, the evolution of attack and defense techniques, and secure development processes. With the application of artificial intelligence and automation technologies, future web security research holds broad prospects in areas such as vulnerability detection and remediation, attack defense, and security engineering. This paper takes the Gruyere platform as its research subject(li2024detecting, ). Through literature review and experimental reproduction, it classifies, analyzes, and validates typical web vulnerabilities while exploring corresponding defense strategies. The research aims to deepen understanding of vulnerability mechanisms and security protection, thereby supporting innovation in web security technology, education and training, and defensive practices(burley2019cybersecurity, ).
@@ 2.1. Web System Security
Within the web security framework, diverse security threats exist at every layer. These threats are intertwined, potentially compromising the entire process from network communication to application logic and data storage. Consequently, they pose serious challenges to the confidentiality, integrity, and availability of the system(hudnall2019educational, ). Distributed Denial-of-Service (DDoS) attacks represent one of the most typical threats at the network layer. Attackers often control a large number of compromised hosts and continuously send automated requests to the target server, thereby exhausting its computational and bandwidth resources. As a result, the server fails to respond to legitimate clients, causing service disruption. Another prevalent threat is Address Resolution Protocol (ARP) spoofing. By constructing forged ARP messages, attackers can map the legitimate IP address of a gateway or host to their own MAC address, thereby conducting man-in-the-middle attacks within a local area network to steal or tamper with data(wei2006preventing, ). At the application layer, SQL injection and Cross-Site Scripting constitute the primary risks. SQL injection exploits insufficient input validation to embed malicious SQL statements into query fields, allowing attackers to bypass authentication, exfiltrate sensitive information, and even implant web shells to gain server control. Cross-Site Scripting, on the other hand, injects malicious JavaScript code into web pages, which is then executed on the client side. This leads to the disclosure of sensitive data, session hijacking, and content manipulation(khochare2011survey, ). In addition, Cross-Site Request Forgery is another common threat, where attackers trick users into unknowingly submitting forged requests, potentially resulting in unauthorized operations such as fund transfers or password changes. Data security challenges are primarily reflected in data exposure and data tampering. Data exposure typically results from application vulnerabilities, configuration errors, or operational mistakes, which can lead to the leakage of credentials, personal information, or business-sensitive data. Data tampering refers to unauthorized modification of database records through various attack vectors, thereby compromising data integrity and adversely affecting business stability and trustworthiness(halfond2006classification, ).
@@ 2.2. Web Security Framework
The core components of a web security framework primarily include network security devices and authentication/encryption mechanisms. Among these, firewalls serve as the primary defense line, typically deployed at the boundary between web servers and external networks. They filter traffic through predefined access control policies, thereby blocking unauthorized connections and data transmission to reduce external intrusion risks(singh2014survey, ). Firewalls can implement granular controls based on parameters such as IP addresses, port numbers, and communication protocols. Intrusion Detection Systems (IDS) monitor network traffic in real-time, issuing immediate alerts upon identifying potential attacks. In contrast, Intrusion Prevention Systems (IPS) possess stronger proactive defense capabilities, automatically taking blocking measures such as discarding malicious packets or restricting abnormal connections(patil2011cross, ). For application-layer security risks, web Application Firewalls (WAFs) serve as specialized security tools, effectively identifying and defending against various web attacks. For identity authentication and access control, authentication mechanisms and authorization management work in tandem to verify user legitimacy while preventing unauthorized access and sensitive information leaks through permission allocation policies. To safeguard data confidentiality and integrity during transmission, common measures include encryption using algorithms like AES and RSA, supplemented by regular data backups for rapid recovery in case of loss or corruption(kieyzun2009automatic, ).
@@ 2.3. Web Security Protection
The framework of web security protection encompasses both technical measures and organizational strategies. From a technical perspective, several key dimensions are involved. First, strict input validation and filtering mechanisms must be enforced to control user-provided data, thereby preventing common threats such as SQL injection and cross-site scripting attacks(patil2015client, ). Second, adherence to secure coding standards is essential, requiring developers and operators to fully understand security specifications, avoid the use of unsafe functions, and apply encryption to sensitive information. In addition, unnecessary system services and network ports should be disabled to minimize the attack surface. At the communication layer, the adoption of secure transmission protocols such as HTTPS ensures the confidentiality and integrity of data during transmission through encryption(shalini2011prevention, ). Furthermore, a comprehensive security auditing and monitoring system should be established to continuously observe and record web application activities, enabling the timely identification of anomalies and potential threats. Audit logs must retain sufficient events and coverage to support traceability and forensic analysis in the event of security incidents.
From a strategic perspective, enhancing security awareness and skills through systematic training and education plays a crucial role. Web developers, system administrators, and end-users should be provided with targeted security training. Developers need to be familiar with common web vulnerabilities and their countermeasures, while administrators should master the configuration and management of security devices and perform regular security assessments and vulnerability scans(mishra2014solving, ). Moreover, a well-structured incident response plan must be developed to enable rapid containment and mitigation of damages when security incidents occur. By integrating technical safeguards with organizational strategies, the overall resilience and robustness of web systems can be significantly strengthened(lu2025movescanner, ).
@@ 3.1. Injection Attack
Injection attacks rank among the most common and severe security threats in web applications. Their fundamental method involves attackers embedding malicious commands within input fields to trick systems into executing unauthorized operations. Such attacks may directly target databases, operating systems, or directory services, leading to data leaks, tampering, or even loss of system control. Depending on the targeted component and exploitation method, injection attacks can be categorized into various types, such as SQL injection, NoSQL injection, command injection, LDAP injection, XPath injection, expression language injection, and ORM injection(shema2010seven, ).
@@ 3.1.1. Database Injection Attacks
SQL injection is one of the most common web application security vulnerabilities. Attackers execute unauthorized database operations by embedding malicious SQL statements within user inputs. Common techniques include Union Query Injection, Error-Based Injection, and Blind Injection. These methods can bypass conventional defenses in various environments, posing severe threats to database confidentiality and integrity(kesharwani2012survey, ). Key defenses against SQL injection include employing prepared statements and parameterized queries to prevent input concatenation, using secure APIs and input validation mechanisms to restrict special characters, and enhancing protection through whitelist strategies. Additionally, adhering to the principle of least privilege by rationally allocating database access rights, supplemented by security measures like web application firewalls, reduces injection attack risks across multiple layers(pandiaraja2015web, ). NoSQL injection has emerged as a significant threat with the increasing adoption of NoSQL databases. The essence of this attack lies in exploiting the flexibility of query languages and syntax features to bypass validation and manipulate data operations. Typical approaches include syntax injection, where malicious payloads are directly embedded into queries, and operator injection, which leverages query operators (e.g., \$ne, \$or, \$where) to achieve unauthorized access. Compared with traditional SQL injection, NoSQL injection demonstrates more diverse forms due to heterogeneous query languages and data models(huang2013craxweb, ).
@@
In practice, attackers often submit malformed or abnormal characters during fuzzing tests and analyze response deviations to identify potential vulnerabilities. For example, in MongoDB-based applications, payloads such as {"username": "admin", "password": {"\$ne": null}} may bypass authentication, while the use of \$where enables arbitrary JavaScript execution, potentially exposing sensitive records. Once exploited, such vulnerabilities may lead to unauthorized data disclosure or a complete compromise of database integrity (rosa2012mitigating, ). Mitigation strategies for NoSQL injection risks can be summarized in several key areas. One approach is to adopt parameterized queries and Object-Relational Mapping (ORM) frameworks such as Mongoose, which eliminate the need for direct string concatenation and thereby reduce the likelihood of injection. Another important measure is the enforcement of strict input validation combined with whitelist mechanisms, which prevents unsafe keywords and special characters from being incorporated into query logic(kong2025uechecker, ). In addition, JavaScript execution within queries should be disabled or restricted to limit the possibility of malicious exploitation. Finally, the principle of least privilege should be applied in access control to minimize the potential impact of an attack. By combining these layers of defense, the resilience of NoSQL databases against injection attacks can be significantly strengthened (owasp2010top, ).
@@ 3.1.2. Command Injection
Command injection is essentially a code execution vulnerability that arises when an application fails to properly validate external inputs, thereby allowing attackers to craft malicious payloads that interfere with application logic and trigger unintended execution of system commands. Once exploited, such vulnerabilities may enable attackers to obtain full control over the server, exfiltrate sensitive data, deploy backdoors, or even conduct lateral movement, posing a severe threat to system security. From a technical perspective, command injection often occurs when the server invokes insecure functions to execute system commands(maffeis2010object, ). Examples include system, exec, passthru, shell_exec, and popen in PHP. System, popen, and subprocess.call in Python, and Runtime.getRuntime().exec() in Java. These functions directly interact with the operating system, and without strict input sanitization, they can be exploited to execute arbitrary commands. Defense strategies against command injection can be grouped into three main aspects(zhang2025penetration, ). The first is rigorous input validation and sanitization, which ensures that malicious characters and command fragments are blocked before they reach the execution environment. The second is the adoption of a whitelist-based restriction mechanism that limits the range of permissible inputs and avoids reliance on high-risk system functions. The third is the use of parameterized execution, which decouples user inputs from system commands and prevents the injection of arbitrary code. When applied together, these measures significantly reduce the likelihood of command injection and enhance the overall security posture of web systems (felmetsger2010toward, ).
@@ 3.1.3. Query Language Injection
The Lightweight Directory Access Protocol, with its distinctive query architecture and syntax rules, is widely used in directory service systems for data retrieval. However, similar to SQL injection, LDAP injection exploits the lack of input validation in applications. By crafting malicious inputs, attackers can manipulate LDAP queries and perform unauthorized operations. Currently, Active Directory Application Mode (ADAM) and OpenLDAP are two commonly adopted implementations in enterprise and network environments(artzi2010finding, ). LDAP injection typically arises in three scenarios. The first scenario involves single-filter injection. When a query contains only one condition, such as (attribute=value), an attacker may insert additional fragments into the input, for example (value)(inject_filter), which results in a concatenated query like (attribute=value)(inject_filter). In OpenLDAP, such attempts are generally ignored, whereas in ADAM, multiple filters are not permitted and the query therefore fails, rendering the injection ineffective. The second scenario concerns multi-filter injection. If the original query employs logical operators such as OR, written as ”|”, attackers can construct inputs such as (|(attribute=value)(inject_filter))(peng2025multicfv, ). In this case, OpenLDAP may parse only the first valid filter, but some LDAP clients or middleware are capable of processing multiple filters simultaneously, which introduces injection risks. The third scenario is syntax-check bypass. Some application frameworks validate filter syntax before sending requests to the server. However, attackers can craft inputs that remain syntactically valid, such as ((value)(injected_filter))(&(1=0)), thereby evading detection and appending unintended filters to the query (garzotto2011enterprise, ).
@@
A common method for detecting LDAP injection is to submit requests that contain intentionally malformed characters and then observe the server’s response. If error messages indicate that the query was executed, the application may be at risk of injection. Defensive strategies can be summarized as follows. First, secure functions such as ldap_escape should be used to properly escape user input in order to prevent structural changes to the query. Second, whitelist-based validation needs to be enforced so that only permissible input is accepted. Finally, parameterized queries should be adopted to avoid concatenating user input directly with LDAP statements. These measures can significantly reduce the security risks posed by LDAP injection (offutt2002quality, ).
@@
XPath injection occurs when an attacker crafts malicious input that is embedded into the XPath expressions of a web application, thereby altering query semantics and causing unintended operations. The underlying principle is similar to SQL injection, where manipulating the query structure can bypass authentication or enable unauthorized access to sensitive data. Common attack methods can be summarized in three categories. The first is logical condition bypass, where attackers introduce always true predicates that cause the query to succeed regardless of the original constraints(zou2025malicious, ). The second is string literal escaping, where attackers modify or terminate string literals to alter the query structure. The third is structural modification, where attackers insert additional predicates or path expressions that expand the query scope or enable access to unauthorized information. Defensive strategies follow the same principles as those for other injection attacks. Parameterized queries or variable binding should be adopted to prevent direct string concatenation. Strict input validation is required together with context-aware encoding of special characters. Data exposure should be minimized by avoiding the direct return of raw query results. In addition, the principle of least privilege must be applied to restrict access to sensitive data and reduce the overall security risks(offutt2002quality, ).

### v2

@@ 1. Introduction
With the rapid advancement of internet technology, web systems have become a critical infrastructure for exchanging information and processing business. However, their complexity has also led to the emergence of numerous security vulnerabilities, making web security a core research focus in cybersecurity. Web security not only concerns data protection and privacy but also directly impacts business continuity and system reliability(Chen et al., 2013). Therefore, conducting web security research holds significant theoretical and practical importance for mitigating malicious attack risks and enhancing system robustness(Zangana, 2024). Gruyere is a web application platform developed by Google specifically for security research. It integrates multiple typical vulnerabilities, covering cross-site scripting (XSS), cross-site request forgery (CSRF), and access control flaws(Douligeris and Mitrokotsa, 2004). The platform is easy to operate and closely mirrors real-world scenarios, providing researchers with an experimental environment for vulnerability analysis and verification while serving as an ideal tool for teaching and training(Altaf et al., 2025). Research based on Gruyere can reveal shortcomings in existing protective mechanisms, drive the development of novel defense technologies, and provide references for the industry in vulnerability protection and risk control(Zivanic et al., 2022).
In recent years, scholars worldwide have made positive progress in the field of web security. However, deficiencies remain in vulnerability detection, the evolution of attack and defense techniques, and secure development processes(Murgia et al., 2024). With the application of artificial intelligence and automation technologies, future web security research holds broad prospects in areas such as vulnerability detection and remediation, attack defense, and security engineering. This paper takes the Gruyere platform as its research subject(Li et al., 2024). Through literature review and experimental reproduction, it classifies, analyzes, and validates typical web vulnerabilities while exploring corresponding defense strategies(Lubis et al., 2024). The research aims to deepen understanding of vulnerability mechanisms and security protection, thereby supporting innovation in web security technology, education and training, and defensive practices(Burley and Lewis Jr, 2019).
@@ 2.1. Web System Security
Within the web security framework, diverse security threats exist at every layer. These threats are intertwined, potentially compromising the entire process from network communication to application logic and data storage. Consequently, they pose serious challenges to the confidentiality, integrity, and availability of the system(Hudnall, 2019). Distributed Denial-of-Service (DDoS) attacks represent one of the most typical threats at the network layer. Attackers often control a large number of compromised hosts and continuously send automated requests to the target server, thereby exhausting its computational and bandwidth resources(Jeberson and others, 2024). As a result, the server fails to respond to legitimate clients, causing service disruption. Another prevalent threat is Address Resolution Protocol (ARP) spoofing. By constructing forged ARP messages, attackers can map the legitimate IP address of a gateway or host to their own MAC address, thereby conducting man-in-the-middle attacks within a local area network to steal or tamper with data(Wei et al., 2006). At the application layer, SQL injection and Cross-Site Scripting constitute the primary risks. SQL injection exploits insufficient input validation to embed malicious SQL statements into query fields, allowing attackers to bypass authentication, exfiltrate sensitive information, and even implant web shells to gain server control(Lin and Dong, 2024). Cross-Site Scripting, on the other hand, injects malicious JavaScript code into web pages, which is then executed on the client side. This leads to the disclosure of sensitive data, session hijacking, and content manipulation(Khochare et al., 2011). In addition, Cross-Site Request Forgery is another common threat, where attackers trick users into unknowingly submitting forged requests, potentially resulting in unauthorized operations such as fund transfers or password changes(Al-Zekri et al., 2024). Data security challenges are primarily reflected in data exposure and data tampering. Data exposure typically results from application vulnerabilities, configuration errors, or operational mistakes, which can lead to the leakage of credentials, personal information, or business-sensitive data(Tehrani et al., 2024). Data tampering refers to unauthorized modification of database records through various attack vectors, thereby compromising data integrity and adversely affecting business stability and trustworthiness(Halfond et al., 2006).
@@ 2.2. Web Security Framework
The core components of a web security framework primarily include network security devices and authentication/encryption mechanisms. Among these, firewalls serve as the primary defense line, typically deployed at the boundary between web servers and external networks(Rey and Labayan, 2024). They filter traffic through predefined access control policies, thereby blocking unauthorized connections and data transmission to reduce external intrusion risks(Singh and Sathappan, 2014). Firewalls can implement granular controls based on parameters such as IP addresses, port numbers, and communication protocols. Intrusion Detection Systems (IDS) monitor network traffic in real-time, issuing immediate alerts upon identifying potential attacks(Salim and Lahcen, 2024). In contrast, Intrusion Prevention Systems (IPS) possess stronger proactive defense capabilities, automatically taking blocking measures such as discarding malicious packets or restricting abnormal connections(Patil et al., 2011). For application-layer security risks, web Application Firewalls (WAFs) serve as specialized security tools, effectively identifying and defending against various web attacks. For identity authentication and access control, authentication mechanisms and authorization management work in tandem to verify user legitimacy while preventing unauthorized access and sensitive information leaks through permission allocation policies. To safeguard data confidentiality and integrity during transmission, common measures include encryption using algorithms like AES and RSA, supplemented by regular data backups for rapid recovery in case of loss or corruption(Kieyzun et al., 2009).
@@ 2.3. Web Security Protection
The framework of web security protection encompasses both technical measures and organizational strategies. From a technical perspective, several key dimensions are involved. First, strict input validation and filtering mechanisms must be enforced to control user-provided data, thereby preventing common threats such as SQL injection and cross-site scripting attacks(Patil and Patil, 2015). Second, adherence to secure coding standards is essential, requiring developers and operators to fully understand security specifications, avoid the use of unsafe functions, and apply encryption to sensitive information. In addition, unnecessary system services and network ports should be disabled to minimize the attack surface. At the communication layer, the adoption of secure transmission protocols such as HTTPS ensures the confidentiality and integrity of data during transmission through encryption(Shalini and Usha, 2011). Furthermore, a comprehensive security auditing and monitoring system should be established to continuously observe and record web application activities, enabling the timely identification of anomalies and potential threats. Audit logs must retain sufficient events and coverage to support traceability and forensic analysis in the event of security incidents.
From a strategic perspective, enhancing security awareness and skills through systematic training and education plays a crucial role. Web developers, system administrators, and end-users should be provided with targeted security training. Developers need to be familiar with common web vulnerabilities and their countermeasures, while administrators should master the configuration and management of security devices and perform regular security assessments and vulnerability scans(Mishra et al., 2014). Moreover, a well-structured incident response plan must be developed to enable rapid containment and mitigation of damages when security incidents occur. By integrating technical safeguards with organizational strategies, the overall resilience and robustness of web systems can be significantly strengthened(Luo et al., 2025).
@@ 3.1. Injection Attack
Injection attacks rank among the most common and severe security threats in web applications. Their fundamental method involves attackers embedding malicious commands within input fields to trick systems into executing unauthorized operations(Paul et al., 2024). Such attacks may directly target databases, operating systems, or directory services, leading to data leaks, tampering, or even loss of system control(Liao et al., 2024). Depending on the targeted component and exploitation method, injection attacks can be categorized into various types, such as SQL injection, NoSQL injection, command injection, LDAP injection, XPath injection, expression language injection, and ORM injection(Shema, 2010).
@@ 3.1.1. Database Injection Attacks
SQL injection is one of the most common web application security vulnerabilities. Attackers execute unauthorized database operations by embedding malicious SQL statements within user inputs(Shi et al., 2024). Common techniques include Union Query Injection, Error-Based Injection, and Blind Injection. These methods can bypass conventional defenses in various environments, posing severe threats to database confidentiality and integrity(Kesharwani and Deshpande, 2012). Key defenses against SQL injection include employing prepared statements and parameterized queries to prevent input concatenation, using secure APIs and input validation mechanisms to restrict special characters, and enhancing protection through whitelist strategies. Additionally, adhering to the principle of least privilege by rationally allocating database access rights, supplemented by security measures like web application firewalls, reduces injection attack risks across multiple layers(Pandiaraja and Manikandan, 2015). NoSQL injection has emerged as a significant threat with the increasing adoption of NoSQL databases. The essence of this attack lies in exploiting the flexibility of query languages and syntax features to bypass validation and manipulate data operations. Typical approaches include syntax injection, where malicious payloads are directly embedded into queries, and operator injection, which leverages query operators (e.g., \$ne, \$or, \$where) to achieve unauthorized access. Compared with traditional SQL injection, NoSQL injection demonstrates more diverse forms due to heterogeneous query languages and data models(Huang et al., 2013).
@@
In practice, attackers often submit malformed or abnormal characters during fuzzing tests and analyze response deviations to identify potential vulnerabilities. For example, in MongoDB-based applications, payloads such as {"username": "admin", "password": {"\$ne": null}} may bypass authentication, while the use of \$where enables arbitrary JavaScript execution, potentially exposing sensitive records. Once exploited, such vulnerabilities may lead to unauthorized data disclosure or a complete compromise of database integrity (Rosa et al., 2012). Mitigation strategies for NoSQL injection risks can be summarized in several key areas. One approach is to adopt parameterized queries and Object-Relational Mapping (ORM) frameworks such as Mongoose, which eliminate the need for direct string concatenation and thereby reduce the likelihood of injection. Another important measure is the enforcement of strict input validation combined with whitelist mechanisms, which prevents unsafe keywords and special characters from being incorporated into query logic(Kong et al., 2025). In addition, JavaScript execution within queries should be disabled or restricted to limit the possibility of malicious exploitation. Finally, the principle of least privilege should be applied in access control to minimize the potential impact of an attack. By combining these layers of defense, the resilience of NoSQL databases against injection attacks can be significantly strengthened (OWASP, 2010).
@@ 3.1.2. Command Injection
Command injection is essentially a code execution vulnerability that arises when an application fails to properly validate external inputs, thereby allowing attackers to craft malicious payloads that interfere with application logic and trigger unintended execution of system commands(Otoom, 2025). Once exploited, such vulnerabilities may enable attackers to obtain full control over the server, exfiltrate sensitive data, deploy backdoors, or even conduct lateral movement, posing a severe threat to system security. From a technical perspective, command injection often occurs when the server invokes insecure functions to execute system commands(Maffeis et al., 2010). Examples include system, exec, passthru, shell_exec, and popen in PHP. System, popen, and subprocess.call in Python, and Runtime.getRuntime().exec() in Java. These functions directly interact with the operating system, and without strict input sanitization, they can be exploited to execute arbitrary commands. Defense strategies against command injection can be grouped into three main aspects(Zhang et al., 2025a). The first is rigorous input validation and sanitization, which ensures that malicious characters and command fragments are blocked before they reach the execution environment(Neef et al., 2024). The second is the adoption of a whitelist-based restriction mechanism that limits the range of permissible inputs and avoids reliance on high-risk system functions(Eberhardt and Milánkovich, 2024). The third is the use of parameterized execution, which decouples user inputs from system commands and prevents the injection of arbitrary code. When applied together, these measures significantly reduce the likelihood of command injection and enhance the overall security posture of web systems (Felmetsger et al., 2010).
@@ 3.1.3. Query Language Injection
The Lightweight Directory Access Protocol, with its distinctive query architecture and syntax rules, is widely used in directory service systems for data retrieval. However, similar to SQL injection, LDAP injection exploits the lack of input validation in applications. By crafting malicious inputs, attackers can manipulate LDAP queries and perform unauthorized operations. Currently, Active Directory Application Mode (ADAM) and OpenLDAP are two commonly adopted implementations in enterprise and network environments(Artzi et al., 2010). LDAP injection typically arises in three scenarios. The first scenario involves single-filter injection. When a query contains only one condition, such as (attribute=value), an attacker may insert additional fragments into the input, for example (value)(inject_filter), which results in a concatenated query like (attribute=value)(inject_filter). In OpenLDAP, such attempts are generally ignored, whereas in ADAM, multiple filters are not permitted and the query therefore fails, rendering the injection ineffective. The second scenario concerns multi-filter injection. If the original query employs logical operators such as OR, written as ”|”, attackers can construct inputs such as (|(attribute=value)(inject_filter))(Peng et al., 2025b). In this case, OpenLDAP may parse only the first valid filter, but some LDAP clients or middleware are capable of processing multiple filters simultaneously, which introduces injection risks. The third scenario is syntax-check bypass. Some application frameworks validate filter syntax before sending requests to the server. However, attackers can craft inputs that remain syntactically valid, such as ((value)(injected_filter))(&(1=0)), thereby evading detection and appending unintended filters to the query (Garzotto, 2011).
@@
A common method for detecting LDAP injection is to submit requests that contain intentionally malformed characters and then observe the server’s response. If error messages indicate that the query was executed, the application may be at risk of injection. Defensive strategies can be summarized as follows. First, secure functions such as ldap_escape should be used to properly escape user input in order to prevent structural changes to the query. Second, whitelist-based validation needs to be enforced so that only permissible input is accepted. Finally, parameterized queries should be adopted to avoid concatenating user input directly with LDAP statements. These measures can significantly reduce the security risks posed by LDAP injection (Offutt, 2002).
@@
XPath injection occurs when an attacker crafts malicious input that is embedded into the XPath expressions of a web application, thereby altering query semantics and causing unintended operations. The underlying principle is similar to SQL injection, where manipulating the query structure can bypass authentication or enable unauthorized access to sensitive data. Common attack methods can be summarized in three categories. The first is logical condition bypass, where attackers introduce always true predicates that cause the query to succeed regardless of the original constraints(Zou et al., 2025). The second is string literal escaping, where attackers modify or terminate string literals to alter the query structure. The third is structural modification, where attackers insert additional predicates or path expressions that expand the query scope or enable access to unauthorized information. Defensive strategies follow the same principles as those for other injection attacks. Parameterized queries or variable binding should be adopted to prevent direct string concatenation. Strict input validation is required together with context-aware encoding of special characters. Data exposure should be minimized by avoiding the direct return of raw query results. In addition, the principle of least privilege must be applied to restrict access to sensitive data and reduce the overall security risks(Offutt, 2002).